16/Sep/99 Evening Shift

Peter Shanahan (shanahan@fnal.gov)
Fri, 17 Sep 1999 00:16:36 -0500

Shift Summary for eve shift, 16/Sep/99

On Shift: Ricardo Barbosa, P. Shanahan

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 2

Accounting of lost time:

Switchover from E832 to E799, continued

========================================================================
Runs ended this shift:
Run 14540 - DBC timing scan
Run 14550 - LASER SCAN

Tapes brought to FCC this shift:
(including tapes for online split)

None

========================================================================
Accesses during shift:

Open Access for most of shift.

========================================================================
Overview of shift:
E.g., Switchover stuff

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15:30 SCALER_CHECK limits for E799 modified (BT)
15:45 ET Chilled water + LV back on (ST,JCH)
16:00 VV' HV turned on. Light leak found + fixed (VJ,SL)
16:15 TRD HV turned on (VJ)
16:50 REG moved to "way out" position. (DJ). NM3 gate
valve opened: ion guage jumps to 80 uTorr. Closed.
17:32 Collimator jaws opened to 2cm x 2cm (BT,EB)
17:35 2 new Ar/Ethane bottles online on east side.
18:05 Crash buttons enabled on NM4AN (BT, DJ)
14540 18:25 End DPMT timing scan (PSh)
18:30 1/2ns delay added to ET RF delay (total=9.5ns) (JCH)
21:00 New DBC control words installed (PSh)
21:10 Vacuum techs leak checked NM2+NM3. No leak found.
They suggest it might be moisture from collimator
insert.
22:15 Search and secure. Gate valve opened (PSh). Ion
gauge jumps to 60uTorr, and decays to 40 uTorr. DJ
says this is OK for beam.
23:59 Vac/cryo problems at E12. At least 1hour for beam.


========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Don't page Jason L. for frowny bananas. He'll fix it tomorrow.
Page Bob T, Sasha L, Arun. when muons are 30 minutes away.